# Don’t be lazy: CompleteP enables compute-efficient deep transformers

### v1

@@ Abstract
We study compute efficiency of LLM training when using different parameterizations, i.e., rules for adjusting model and optimizer hyperparameters (HPs) as model size changes. Some parameterizations fail to transfer optimal base HPs (such as learning rate) across changes in model depth, requiring practitioners to either re-tune these HPs as they scale up (expensive), or accept sub-optimal training when re-tuning is prohibitive. Even when they achieve HP transfer, we develop theory to show parameterizations may still exist in the lazy learning regime where layers learn only features close to their linearization, preventing effective use of depth and nonlinearity. Finally, we identify and adopt the unique parameterization we call CompleteP that achieves both depth-wise HP transfer and non-lazy learning in all layers. CompleteP enables a wider range of model width/depth ratios to remain compute-efficient, unlocking shapes better suited for different hardware settings and operational contexts. Moreover, CompleteP enables 12-34% compute efficiency improvements over the prior state-of-the-art.
@@ 1 Introduction
where $\mathcal{F}_{\ell}$ is the $\ell$’th residual block (for transformers, these are MLP and attention blocks). The depth-dependent re-scaling factor is governed by a single parameter $\alpha\in[0.5,1]$, where a larger $\alpha$ corresponds to stronger downscaling. Setting $\alpha=1$ gives a parameterization we call CompleteP because it is the unique parameterization ensuring complete feature learning as width and depth are scaled (Section 6). Since $\alpha=0.5$ and $\alpha=1$ are the two most promising potential candidates for depth scaling [6, 7, 8], we compare these two values and find $\alpha=1$ is consistently more compute-efficient through better HP transfer and faster pre-training (Figure 1). To realize these gains, however, we must extend the parameterization to include principled re-scalings of LayerNorm (LN) and bias learning rates, AdamW’s weight decay $\lambda$, and AdamW’s $\epsilon$ (Table 1). CompleteP is simple to implement, yet yields superior upstream and downstream performance, with gains over alternative approaches increasing with model depth. To summarize, our contributions are:
@@
  • Optimal shape and compute efficiency. For SP, µP, and $\alpha=\{0.5,1\}$, we construct scaling laws over models ranging from $75$M to $1.9$B total parameters, trained at a compute-optimal frontier of 20 tokens per parameter (TPP) [3]. We use this to revisit the question of compute efficient transformer shapes (i.e. width-to-depth ratio $N\!:\!L$). The 1.9B parameter models (i.e. with 1.5B non-embedding parameters) trained with $\alpha=1$ achieve 11.8% and 34.4% FLOP savings over µP for optimally-shaped and 179-layer models, respectively (Figure 4).
@@ 2 Related work
Early theoretical approaches for choosing HPs analyzed networks at initialization, seeking to guarantee different notions of numerical stability of the first forward and backward pass [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Subsequent work devised two parameterizations with consistent training dynamics at infinite width: the Neural Tangent Kernel (NTK) parameterization in which the model converges to its linearization around initialization [29, 30, 31] and the mean-field/µP parameterization [4, 6, 8, 10, 32, 33, 34, 35, 36, 37, 38, 39]. Training at the simultaneous limits of width and depth was explored in [40] for fully connected networks, in [6, 7] for vanilla ResNets and in [8] for transformers. Limiting descriptions of training in the mean-field/µP parameterization are difficult to study analytically but give a clear proposal for HP transfer by requiring consistent dynamics of hidden layer representations across model scale [41, 42, 43, 44, 32].
@@
Empirical approaches to HP transfer in [49, 50] normalize layer outputs and learning rates to yield HP transfer across width, consistent with the µP prescription. HP transfer across depth is often more ad hoc: [49] proposes something resembling $\alpha=0.5$, while Large et al. [50] proposes $\mathbf{h}^{\ell+1}={(\frac{L-1}{L}\cdot\mathbf{I}+\frac{1}{L}\cdot\mathcal{F}_{\ell}(\mathbf{h}^{\ell})}^{L}$, reminiscent of setting $\alpha=1$. Several other works, though not directly focused on HP transfer, aim to stabilize training in deep ResNets through LN modifications. Sun et al. [51] advocates multiplying pre-LN output in layer $\ell$ by $\ell^{-0.5}$, similar to taking $\alpha=0.5$. Similarly, [52] applies LN to both the input and output of each residual block, avoiding exponential-in-depth activation and gradient variance. Finally, [53] adjusts the scale of weights in each residual block, up-weights residual branch contributions, and inserts LN after the residual add. While this allows one to train very deep networks, it does not achieve depth HP transfer.
@@
Theoretical and empirical work have studied the optimal transformer shape. Notably Kaplan et al. [2], using SP and large-scale tests, showed a wide range of $N\!:\!L$ were close to compute-optimality, with $N\!:\!L\approx 100$ being the optimal ratio. This finding guides modern LLM shapes, which often fix $N\!:\!L\approx 100$ when parameters and tokens are scaled [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Yet SP does not fairly admit stable width and depth scaling, which undermines the prior conclusion. In Section 5 we revisit the transformer $N\!:\!L$ study with proper width and depth scaling control with CompleteP and show that even $N\!:\!L\approx 10$ remains close to compute optimality. Rather than fixing $N\!:\!L$, other works perform empirical searches to estimate scaling exponents for $N$ and $L$ [73, 74]. Levine et al. [75] propose a theory for transformer capacity based on separation rank, predict that optimal $N\!:\!L$ increases with parameter count, and provide empirical validation. Mixture of experts (MoE) [76] enable transformer capacity to scale without increasing depth. Similarly, [77] propose parallel sub-networks and argue that large depth isn’t necessary for competitive performance. We focus on dense transformers and consider MoE and parallel sub-networks as out of scope.
@@ 3 Methodology
We train decoder-only transformer language models [78], with pre-normalization, on the SlimPajama dataset [79]. We refer the reader to Appendix I for full methodology of all experiments. Table 1 provides an overview of how to implement the parameterizations we test in this paper. [^1] For the embedding layer, we choose an initialization variance scale independent of input dimension due to the input data being one-hot encoded. If the input data is dense, then we would require a pre-factor of $N_{\mathrm{in}}^{-1/2}$. We consider models with depth $L$ (or $2L$ residual blocks to account for both MLP and attention blocks), and width $N$ (e.g., residual activations $h^{\ell}\in\mathbb{R}^{N}$). We define adjustments in terms of width multiplier $m_{N}=N/N_{\text{base}}$ and depth multiplier $m_{L}=L/L_{\text{base}}$, where $N_{\text{base}},L_{\text{base}}$ are the width and the depth of a base model. For the base model, i.e., when $m_{N}=1,m_{L}=1$, all parameterizations are equivalent. We use $N_{\text{base}}=256,L_{\text{base}}=2$. In experiments where only $m_{N}=1$, SP is equivalent to µP; such results are labelled “SP/µP”. We extend the infinite depth parameterizations in the literature [6, 7, 8] with corrections for bias learning rate $\eta$, LayerNorm $\eta$, AdamW $\epsilon$, and weight decay $\lambda$. See Appendix E for derivations of these extensions and Appendix F for empirical verification of $\epsilon$ scaling. We provide a minimal implementation of $\alpha\in\{0.5,1.0\}$ at: https://github.com/EleutherAI/nanoGPT-mup/tree/completep.
@@
Embed Init. Var. $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$
Embed LR (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$
@@
Final-LN LR WandB (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}\color[rgb]{.132,.545,.132}\definecolor[named]{pgfstrokecolor}{rgb}{.132,.545,.132}{m_{L}}^{\alpha-1}$
Unembed Init. Var. $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$
Unembed LR (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$
Unembed Fwd. $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}$ $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}m_{N}^{-1}$ $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}m_{N}^{-1}$
@@ Results
Finding 5: As model scale increases, CompleteP ($\alpha=1$) uniquely enables deep-narrow models (small $N\!:\!L$) to remain close to compute-optimal.
@@ 6 Desiderata for Hyperparameter Transfer
This desideratum prescribes how to scale weight variances (or pre-factors) as in Table 1. It also constrains the value of $\alpha$ to be at least $0.5$ (see Section D.1). It can be viewed as a numerical stability condition and has been-well studied in the signal-propagation literature [11, 16, 17, 19, 20, 21, 40]. We review why this desideratum imposes $\alpha\geq 1/2$ in Section D.1.
@@
The network parameterization satisfies complete feature learning, i.e. neither the hidden layers $\{h^{\ell}\}_{\ell\in[L]}$ nor the model output $f$ are are lazy with respect to any subset of model parameters.
@@
Figure 6: Only $\alpha=1$ achieves complete feature learning with stability. SP/$\mu$P diverges at large depth (missing points are NaNs), while $\alpha=0.5$ converges to the linearization at a rate of $1/\sqrt{L}$. The y axis is the left hand side of Eq. 4.
@@
Given that $\alpha=1$ is the unique parameterization that ensures complete feature learning while scaling both width and depth, we dub it CompleteP.
@@ 7 Conclusion
To understand the salutary effects of CompleteP, we formalized a set of desiderata for designing parameterizations. They include variants of desiderata from prior work as well as a novel property, which we called complete feature learning, that distinguishes CompleteP. This parameterization is simple and light weight to implement, and we released a public code base to facilitate reproduction and further study. Our empirical tests were conducted on models with up to 1.5B non-embedding parameters. In future work, we plan to train large, state-of-the-art LLMs that leverage CompleteP’s utility in HP transfer and compute-efficient training.

### v2

@@ Abstract
We study compute efficiency of LLM training when using different parameterizations, i.e., rules for adjusting model and optimizer hyperparameters (HPs) as model size changes. Some parameterizations fail to transfer optimal base HPs (such as learning rate) across changes in model depth, requiring practitioners to either re-tune these HPs as they scale up (expensive), or accept sub-optimal training when re-tuning is prohibitive. Even when they achieve HP transfer, we develop theory to show parameterizations may still exist in the lazy learning regime where layers learn only features close to their linearization, preventing effective use of depth and nonlinearity. Finally, we identify and adopt the parameterization we call CompleteP that achieves both depth-wise HP transfer and non-lazy learning in all layers. CompleteP enables a wider range of model width/depth ratios to remain compute-efficient, unlocking shapes better suited for different hardware settings and operational contexts. Moreover, CompleteP enables 12-34% compute efficiency improvements over the prior state-of-the-art.
@@ 1 Introduction
where $\mathcal{F}_{\ell}$ is the $\ell$’th residual block (for transformers, these are MLP and attention blocks). The depth-dependent re-scaling factor is governed by a single parameter $\alpha\in[0.5,1]$, where a larger $\alpha$ corresponds to stronger downscaling. Setting $\alpha=1$ gives a parameterization we call CompleteP because it is the unique $\alpha$ ensuring complete feature learning as width and depth are scaled (Section 6). Since $\alpha=0.5$ and $\alpha=1$ are the two most promising potential candidates for depth scaling [6, 7, 8], we compare these two values and find $\alpha=1$ is consistently more compute-efficient through better HP transfer and faster pre-training (Figure 1). To realize these gains, however, we must extend the parameterization to include principled re-scalings of LayerNorm (LN) and bias learning rates, AdamW’s weight decay $\lambda$, and AdamW’s $\epsilon$ (Table 1). CompleteP is simple to implement, yet yields superior upstream and downstream performance, with gains over alternative approaches increasing with model depth. To summarize, our contributions are:
@@
  • Optimal shape and compute efficiency. For SP, µP, and $\alpha=\{0.5,1\}$, we construct scaling laws over models ranging from $75$M to $1.9$B total parameters, trained at a compute-optimal frontier of 20 tokens per parameter (TPP) [3]. We use this to revisit the question of compute-efficient transformer shapes (i.e. width-to-depth ratio $N\!:\!L$). The 1.9B parameter models (i.e. with 1.5B non-embedding parameters) trained with $\alpha=1$ achieve 11.8% and 34.4% FLOP savings over µP for optimally-shaped and 179-layer models, respectively (Figure 4).
@@ 2 Related work
Early theoretical approaches for choosing HPs analyzed networks at initialization, seeking to guarantee different notions of numerical stability of the first forward and backward pass [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Subsequent work devised two parameterizations with consistent training dynamics at infinite width: the Neural Tangent Kernel (NTK) parameterization in which the model converges to its linearization around initialization [29, 30, 31] and the mean-field/µP parameterization [4, 6, 8, 10, 32, 33, 34, 35, 36, 37, 38, 39]. Training at the simultaneous limits of width and depth was explored in [40] for fully connected networks, in [6, 7] for vanilla ResNets and in [8] for transformers. Infinite-limit descriptions of training in the mean-field/µP parameterization are difficult to study analytically but give a clear proposal for HP transfer by requiring consistent dynamics of hidden layer representations across model scale [41, 42, 43, 44, 32].
@@
Empirical approaches to HP transfer in [49, 50] normalize layer outputs and learning rates to yield HP transfer across width, consistent with the µP prescription. HP transfer across depth is often more ad hoc: [49] proposes something resembling $\alpha=0.5$, while Large et al. [50] proposes $\mathbf{h}^{\ell+1}={(\frac{L-1}{L}\cdot\mathbf{I}+L^{-1}\cdot\mathcal{F}_{\ell}(\mathbf{h}^{\ell}))}^{L}$, reminiscent of setting $\alpha=1$. Several other works, though not directly focused on HP transfer, aim to stabilize training in deep ResNets through LN modifications. Sun et al. [51] advocates multiplying pre-LN output in layer $\ell$ by $\ell^{-0.5}$, similar to taking $\alpha=0.5$. Similarly, [52] applies LN to both the input and output of each residual block, avoiding exponential-in-depth activation and gradient variance. Finally, [53] adjusts the scale of weights in each residual block, up-weights residual branch contributions, and inserts LN after the residual add. While this allows one to train very deep networks, it does not achieve depth HP transfer.
@@
Theoretical and empirical work have studied the optimal transformer shape. Notably Kaplan et al. [2], using SP and large-scale tests, showed a wide range of $N\!:\!L$ were close to compute-optimality, with $N\!:\!L\approx 100$ being the optimal ratio. This finding guides modern LLM shapes, which often fix $N\!:\!L\approx 100$ when parameters and tokens are scaled [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Yet SP does not fairly admit stable width and depth scaling, which undermines the prior conclusion. McLeish et al. [73] adopt a $\eta=\eta_{\text{base}}/\sqrt{L}$ parameterization resembling incomplete $\alpha=0.5$ and study compute-optimal $N\!:\!L$. However, their parameterization does not admit training stability (Figure 7) or HP transfer so their deeper models were disadvantaged, causing them to conclude shallower models are optimal. In Section 5 we revisit the transformer $N\!:\!L$ study with proper width and depth scaling control with CompleteP and show that even $N\!:\!L\approx 10$ remains close to compute optimality. Rather than fixing $N\!:\!L$, other works perform empirical searches to estimate scaling exponents for $N$ and $L$ [74, 75]. Levine et al. [76] propose a theory for transformer capacity based on separation rank, predict that optimal $N\!:\!L$ increases with parameter count, and provide empirical validation. Mixture of experts (MoE) [77] enable transformer capacity to scale without increasing depth. Similarly, [78] propose parallel sub-networks and argue that large depth isn’t necessary for competitive performance. We focus on dense transformers and consider MoE and parallel sub-networks as out of scope.
@@ 3 Methodology
We train decoder-only transformer language models [79], with pre-normalization, on the SlimPajama dataset [80]. We refer the reader to Appendix I for full methodology of all experiments. Table 1 provides an overview of how to implement the parameterizations we test in this paper. [^1] For the embedding layer, we choose an initialization variance scale independent of input dimension due to the input data being one-hot encoded. If the input data is dense, then we would require a pre-factor of $N_{\mathrm{in}}^{-1/2}$. We consider models with depth $L$ (or $2L$ residual blocks to account for both MLP and attention blocks), and width $N$ (e.g., residual activations $h^{\ell}\in\mathbb{R}^{N}$). We define adjustments in terms of width multiplier $m_{N}=N/N_{\text{base}}$ and depth multiplier $m_{L}=L/L_{\text{base}}$, where $N_{\text{base}},L_{\text{base}}$ are the width and the depth of a base model. For the base model, i.e., when $m_{N}=1,m_{L}=1$, all parameterizations are equivalent. We use $N_{\text{base}}=256,L_{\text{base}}=2$. In experiments where only $m_{N}=1$, SP is equivalent to µP; such results are labelled “SP/µP”. We extend the infinite depth parameterizations in the literature [6, 7, 8] with corrections for bias learning rate $\eta$, LayerNorm $\eta$, AdamW $\epsilon$, and weight decay $\lambda$. See Appendix E for derivations of these extensions and Appendix F for empirical verification of $\epsilon$ scaling [^2] In practice, we observe that a “small enough” $\epsilon$ is sufficient to achieve consistent dynamics and transfer across width depth. See Appendix I for details.. We provide a minimal implementation of $\alpha\in\{0.5,1.0\}$ at: https://github.com/EleutherAI/nanoGPT-mup/tree/completep.
@@
Emb. Init. Var. $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$
Emb. LR (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$
@@
Final-LN LR (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$
Unemb. Init. Var. $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma_{\text{base}}^{2}$
Unemb. LR (AdamW) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\eta_{\text{base}}$
Unemb. Fwd. $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}$ $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}m_{N}^{-1}$ $\mathbf{X}^{L}\mathbf{W}^{\top}_{\text{unemb}}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}m_{N}^{-1}$
@@
AdamW $\epsilon$ (Emb. & Unemb.) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon_{\text{base}}$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon_{\text{base}}$
@@ Results
Finding 5: As model scale increases, CompleteP ($\alpha=1$) enables deep-narrow models (small $N\!:\!L$) to remain close to compute-optimal.
@@ 6 Desiderata for Hyperparameter Transfer
This desideratum prescribes how to scale weight variances (or pre-factors) as in Table 1. It also constrains the value of $\alpha$ to be at least $0.5$ (see Section D.1). It can be viewed as a numerical stability condition and has been well-studied in the signal-propagation literature [11, 16, 17, 19, 20, 21, 40]. We review why this desideratum imposes $\alpha\geq 1/2$ in Section D.1.
@@
The network parameterization satisfies complete feature learning, i.e. neither the hidden layers $\{h^{\ell}\}_{\ell\in[L]}$ nor the model output $f$ are lazy with respect to any subset of model parameters.
@@
Figure 6: Only $\alpha=1$ achieves complete feature learning with stability. SP/$\mu$P diverges at large depth (triangles are NaNs), while $\alpha=0.5$ converges to the linearization at a rate of $1/\sqrt{L}$. The y axis is the left hand side of Eq. 4.
@@
Given that $\alpha=1$ is the unique $\alpha$ that ensures complete feature learning while scaling both width and depth, we dub it CompleteP.
@@ 7 Conclusion
To understand the salutary effects of CompleteP, we formalized a set of desiderata for designing parameterizations. They include variants of desiderata from prior work as well as a novel property, which we called complete feature learning, that distinguishes CompleteP. This parameterization is simple to implement, and we released a public code base to facilitate reproduction and further study. Our empirical tests were conducted on models with up to 1.5B non-embedding parameters. In future work, we plan to train large, state-of-the-art LLMs that leverage CompleteP’s utility in HP transfer and compute-efficient training.
Acknowledgments and Disclosure of Funding
BB and LN thank the Google PhD fellowship for support. BH is supported by a Sloan Fellowship, NSF DMS-2143754, and NSF DMS-2133806. BB and C.P. acknowledge support by NSF grant DMS-2134157, NSF CAREER Award IIS-2239780, DARPA grant DIAL-FP-038, a Sloan Research Fellowship, and The William F. Milton Fund from Harvard University. C.P. and B.B’s work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.